2x+3y=40 and 2x+2y=20?
Answer by chitra(359)

The given equations are:

2x + 3y = 40 -----------(1)

2x + 2y = 20 -----------(2)

Subtracting EQN (1) from (2), we get:

y = 20

Substituing this in any one of the equations, we get:

2x + 2(20) = 20

2x + 40 = 20

2x = 20 - 40

2x = -20

x = - 20/2

x = -10

Hence, the values of x = -10 and y = 20 satisfies both of the equations.

Hence, the solution.